OpenStudy (anonymous):

Which equation is in slope-intercept form and represents a line with slope 3 through the point (9, –4)?

OpenStudy (dessyj1):

(y-y1)=m(x-x1) is your formula

OpenStudy (dessyj1):

sub in your y-value for y1 and sub in your x-value for x1

OpenStudy (dessyj1):

(y+4)=3(x-9) (y+4)=3x-27 y=3x-31 you should get y=3x-31

OpenStudy (triciaal):

welcome slope intercept form y = mx + b m is the slope change in y / change in x y2 - y1 / x2 - x1 for 2 points on the line (x1, y1) and (x2, y2) b is the y-intercept the value of y when x = 0

OpenStudy (triciaal):